# Erdélyi-type integrals for $F_{K}$ function and their $q$-analogues

### v1

@@ Abstract
In this paper, we revisit the recent result of Luo, Xu, and Raina [Fractal Frac. 6 (3) (2022)] on an Erdélyi-type integral for Saran’s three-variable hypergeometric function $F_{K}$. We provide a new proof of this integral and derive an attractive new integral related to Appell’s function $F_{2}$. A further extension on the $L$-variable $F_{K}$ function, which appears in physics, is also discussed. Furthermore, we prove various $q$-Erdélyi-type integrals for the $q$-analogue of the $F_{K}$-function. An interesting discrete analogue is also included. We also provide a valuable compilation of the sources for known Erdélyi-type integrals of many different hypergeometric functions in the Appendix.
@@
E-mail: mathwinnie@live.com, mathwinnie@dhu.edu.cn
@@
E-mail: guoliangjia77@outlook.com
@@ 1.1 Background and motivation
where $\alpha,\beta\in\mathbb{C}$ and $\gamma\notin\mathbb{Z}_{\leq 0}:=\{0,-1,-2,\cdots\}$ (see [13, p. 56] and [38, p. 384]). It was shown by Euler that [13, p. 59, Eq. (10)]
@@
Here and throughout the paper, we let $\mu_{\alpha,\beta}$ be the Dirichlet measure defined by [6, p. 64, Definition 4.4-1]
@@
$\int_{0}^{1}\int_{0}^{1}(1-zts)^{-\gamma}\mathrm{d}\mu_{\beta,\gamma-\beta}(t)\mathrm{d}\mu_{\alpha,\gamma-\alpha}(s),$
which is symmetric in $\alpha$ and $\beta$. If we integrate with respect to $t$ first, then we obtain the second integral (2).
@@
In (7), $\mu_{\alpha,\beta,\gamma,\eta}$ is the hypergeometric measure defined by [32, p. 6, Eq. (11)]
@@
It is particularly noteworthy that Erdélyi’s integrals are not only of computational interest but also have many important and profound applications. Gasper [17] has shown that (6) provides a natural way to the formulas of Dirichlet-Mehler type for the Jacobi polynomials and the generalized Legendre functions. Sprinkhuizen-Kuyper [47] used (6) to prove a compositional property for her fractional integral operator $I_{\nu}^{\mu,\lambda}$. The integral (6) also plays an important role in the proof of product formula for biangle polynomials [25]. The proofs of the compositional properties of Saigo’s fractional integral operators $I^{\alpha,\beta,\eta}$ and $J^{\alpha,\beta,\eta}$ are based on integrals (5) and (6) [44]. Raina [39] showed that the integral (5) can be used to construct a solution to a certain Abel-type integral equation involving the Appell hypergeometric function $F_{3}$ in the kernel.
We refer to integrals that have a form similar to integrals (5)–(7) as Erdélyi-type integrals. More precisely, for a given hypergeometric function, its Erdélyi-type integrals can generally express it as an integral of the product of two functions of the same type (see [32, p. 2]).
@@ 1.2 The plan of the work
However, the initial proof of integral (1.1) was relatively difficult. In fact, Luo et al. first established a very general integral involving the Srivastava-Daoust function through fractional integration by parts, and then reduced it to (1.1) by specializing the parameters. In this paper, we shall provide a direct and simple proof of Theorem 1.1 in Section 3.1. Then in Section 3.2, we shall derive a curious integral representation for Appell function $F_{2}$ from Theorem 1.1. A further extension, which is related to Belitsky’s generalization of $F_{K}$ [4, p. 69], is considered in Section 3.3.
@@
The multiple $q$-shifted factorials are defined by $(a_{1},\cdots,a_{m};q)_{n}=(a_{1};q)_{n}\cdots(a_{m};q)_{n}$, where $n$ could be an integer or $\infty$. Let $f:\mathbb{C}^{k}\rightarrow\mathbb{C}$ be a function. The $k$-dimensional $q$-integral can be defined by (see, for example, [23, p. 1] and [51, p. 241])
@@
The $q$-gamma function $\Gamma_{q}(x)$ is defined by [19, p. 20, Eq. (1.10.1)]
@@
and the $q$-beta function is defined by [19, p. 22, Eq. (1.10.13)]
@@
Note that $\lim_{q\rightarrow 1^{-}}\Gamma_{q}(x)=\Gamma(x)$ and $\lim_{q\rightarrow 1^{-}}B_{q}(x,y)=B(x,y)$. The $q$-integral representation for $B_{q}(x,y)$ is given by [19, p. 23, Eq. (1.11.7)]
@@
where $|x|<1$, $|y|<1$, $|z|<1$ and $_{r}\phi_{s}$ is the basic hypergeometric series defined by [19, p. 4, Eq. (1.2.22)]
@@
The only $q$-integral representation (of Bateman type) for $\Phi_{K}$ is given by Ernst [15, p. 10, Theorem 5]:
@@ 2 Preliminaries
Gasper [18] discussed the $q$-analogues of Erdélyi’s integrals (5)–(7) and provided their discrete analogues. The integral (5) has the following $q$-analogue [18, p. 4, Eq. (1.13)]
@@
We have $\int_{0}^{1}\mathrm{d}\mu_{\eta-\lambda,\gamma-\lambda,\gamma-\lambda+\eta-\nu,\nu}(t;q)=1$ and $\lim_{q\rightarrow 1^{-}}\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t;q)=\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t)$. Then the $q$-analogue of (7) is given by [18, p. 4, Eq. (1.14)]
@@
Finally, we also need the following $q$-hypergeometric function of three variables (see [9, p. 96] and [40, p. 35]):
@@ 3.2 A curious integral representation for $F_{2}$
By using (33) and the Pfaff transformation [48, p. 300, Eq. (79)], we have
@@
We can prove (3.2) directly by using the method described in Section 3.1. In addition, the integral (3.2) can also be derived from Manocha’s integral [34, p. 239, Eq. (13)]:
@@
Another significant example originates in physics. In 2018, Belitsky [4, p. 69] generalized Saran’s $F_{K}$-function of three variables to $L$ variables, i.e.,
@@ 4.1 A Joshi-Vyas type theorem
Inspired by the formula [21, p. 133] of Joshi and Vyas and its useful $q$-analogue [22, p. 646], we establish–using the measure in (19)–the following theorem of Joshi-Vyas type.
@@
When $k=1$, the formula (4.1) reduces to the result obtained by Joshi and Vyas. It can also be considered as Ernst’s $q$-analogue for Exton’s result (see [14, p. 189, Theorem 3.1]).
@@
This integral (4.3) is clearly a $q$-analogue of the integral by Luo et al. [32, Theorem 1]. In particular, when $x=0$, (4.3) reduces to
@@
which is the $q$-analogue of Koschmieder’s integral representation for $F_{2}$ ([32, Corollary 1]). Here, $\Phi_{2}$ and $\Phi_{0:2;2}^{1:2;2}$ denote the $q$-Appell function [19, p. 283, Eq. (10.2.6)] and the $q$-Kampé de Fériet function [48, p. 349], respectively.
@@
This integral is the $q$-analogue of the integral by Luo and Raina [31, p. 14, Theorem 4.1].
@@
Let $f:\mathbb{C}\rightarrow\mathbb{C}$ be a function of one variable. We define the $q$-shift operator $\mathbb{E}_{q}$ by $\mathbb{E}_{q}f(x)=f(qx)$ $(0<q<1)$ [29, Definition 1.3]. Inductively, for any $n\in\mathbb{Z}_{\geq 1}$,
@@
where $\left[\begin{matrix}k\\ p\end{matrix}\right]_{q}$ denotes the $q$-binomial coefficient [19, p. 24]. Then the use of the $q$-binomial theorem [19, p. 25] gives
@@ 5 Appendix
No funding was received for conducting this study.
Conflicts of interest
The authors declare that there is no conflict of interest.

### v2

@@ Abstract
In this paper, we revisit the recent result of Luo, Xu, and Raina [Fractal Fract. 6 (3) (2022)] on an Erdélyi-type integral for Saran’s three-variable hypergeometric function $F_{K}$. We provide a new proof of this integral and derive an attractive new integral related to Appell’s function $F_{2}$. A further extension on the $L$-variable $F_{K}$ function, which appears in physics, is also discussed. Furthermore, we prove various $q$-Erdélyi-type integrals for the $q$-analogue of the $F_{K}$-function. An interesting discrete analogue is also included. We also provide a valuable compilation of the sources for known Erdélyi-type integrals of many different hypergeometric functions in the Appendix.
@@
E-mail: guoliangjia77@outlook.com
@@
E-mail: mathwinnie@live.com, mathwinnie@dhu.edu.cn
@@ 1.1 Background and motivation
where $\alpha,\beta\in\mathbb{C}$ and $\gamma\notin\mathbb{Z}_{\leq 0}:=\{0,-1,-2,\cdots\}$ (see [14, p. 56] and [39, p. 384]). It was shown by Euler that [14, p. 59, Eq. (10)]
@@
Here and throughout the paper, we let $\mu_{\alpha,\beta}$ be the Dirichlet measure defined by [7, p. 64, Definition 4.4-1]
@@
$\int_{0}^{1}\int_{0}^{1}(1-zts)^{-\gamma}\mathrm{d}\mu_{\beta,\gamma-\beta}(t)\mathrm{d}\mu_{\alpha,\gamma-\alpha}(s),$ (4)
which is symmetric in $\alpha$ and $\beta$. If we integrate with respect to $t$ first, then we obtain the second integral (2). Note that after introducing the new variables $x_{1}=s/(1-s)$ and $x_{2}=t/(1-t)$ and making the substitution $z\rightarrow 1-z$ in Erdélyi’s formula (4), we obtain the double integral
$\int_{0}^{\infty}\int_{0}^{\infty}\frac{x_{1}^{\alpha}x_{2}^{\beta}}{(1+x_{1}+x_{2}+zx_{1}x_{2})^{\gamma}}\frac{\mathrm{d}x_{1}\mathrm{d}x_{2}}{x_{1}x_{2}}=\frac{\Gamma(\beta)\Gamma(\gamma-\beta)\Gamma(\alpha)\Gamma(\gamma-\alpha)}{(\Gamma(\gamma))^{2}}{}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};1-z\right].$ (5)
The integral on the left-hand side of (5) is also known as the Euler-Mellin integral (see [3] and [50]).
@@
In (9), $\mu_{\alpha,\beta,\gamma,\eta}$ is the hypergeometric measure defined by [33, p. 6, Eq. (11)]
@@
It is particularly noteworthy that Erdélyi’s integrals are not only of computational interest but also have many important and profound applications. Gasper [18] has shown that (8) provides a natural way to the formulas of Dirichlet-Mehler type for the Jacobi polynomials and the generalized Legendre functions. Virchenko and Fedotova also frequently used the integrals (7) and (8) in their study of the generalized Legendre functions (see [51], [52] and [53]). Sprinkhuizen-Kuyper [48] used (8) to prove a compositional property for her fractional integral operator $I_{\nu}^{\mu,\lambda}$. The integral (8) also plays an important role in the proof of product formula for biangle polynomials [26]. The proofs of the compositional properties of Saigo’s fractional integral operators $I^{\alpha,\beta,\eta}$ and $J^{\alpha,\beta,\eta}$ are based on integrals (7) and (8) [45]. Raina [40] showed that the integral (7) can be used to construct a solution to a certain Abel-type integral equation involving the Appell hypergeometric function $F_{3}$ in the kernel.
We refer to integrals that have a form similar to integrals (7)–(9) as Erdélyi-type integrals. More precisely, for a given hypergeometric function, its Erdélyi-type integrals can generally express it as an integral of the product of two functions of the same type (see [33, p. 2]).
@@ 1.2 The plan of the work
However, the initial proof of integral (1.1) was relatively difficult. In fact, Luo et al. first established a very general integral involving the Srivastava-Daoust function through fractional integration by parts, and then reduced it to (1.1) by specializing the parameters. In this paper, we shall provide a direct and simple proof of Theorem 1.1 in Section 3.1. Then in Section 3.2, we shall derive a curious integral representation for Appell function $F_{2}$ from Theorem 1.1. A further extension, which is related to Belitsky’s generalization of $F_{K}$ [5, p. 69], is considered in Section 3.3.
@@
The multiple $q$-shifted factorials are defined by $(a_{1},\cdots,a_{m};q)_{n}=(a_{1};q)_{n}\cdots(a_{m};q)_{n}$, where $n$ could be an integer or $\infty$. Let $f:\mathbb{C}^{k}\rightarrow\mathbb{C}$ be a function. The $k$-dimensional $q$-integral can be defined by (see, for example, [24, p. 1] and [59, p. 241])
@@
The $q$-gamma function $\Gamma_{q}(x)$ is defined by [20, p. 20, Eq. (1.10.1)]
@@
and the $q$-beta function is defined by [20, p. 22, Eq. (1.10.13)]
@@
Note that $\lim_{q\rightarrow 1^{-}}\Gamma_{q}(x)=\Gamma(x)$ and $\lim_{q\rightarrow 1^{-}}B_{q}(x,y)=B(x,y)$. The $q$-integral representation for $B_{q}(x,y)$ is given by [20, p. 23, Eq. (1.11.7)]
@@
where $|x|<1$, $|y|<1$, $|z|<1$ and $_{r}\phi_{s}$ is the basic hypergeometric series defined by [20, p. 4, Eq. (1.2.22)]
@@
The only $q$-integral representation (of Bateman type) for $\Phi_{K}$ is given by Ernst [16, p. 10, Theorem 5]:
@@ 2 Preliminaries
Gasper [19] discussed the $q$-analogues of Erdélyi’s integrals (7)–(9) and provided their discrete analogues. The integral (7) has the following $q$-analogue [19, p. 4, Eq. (1.13)]
@@
We have $\int_{0}^{1}\mathrm{d}\mu_{\eta-\lambda,\gamma-\lambda,\gamma-\lambda+\eta-\nu,\nu}(t;q)=1$ and $\lim_{q\rightarrow 1^{-}}\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t;q)=\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t)$. Then the $q$-analogue of (9) is given by [19, p. 4, Eq. (1.14)]
@@
Finally, we also need the following $q$-hypergeometric function of three variables (see [10, p. 96] and [41, p. 35]):
@@ 3.2 A curious integral representation for $F_{2}$
By using (35) and the Pfaff transformation [49, p. 300, Eq. (79)], we have
@@
We can prove (3.2) directly by using the method described in Section 3.1. In addition, the integral (3.2) can also be derived from Manocha’s integral [35, p. 239, Eq. (13)]:
@@
Another significant example originates in physics. In 2018, Belitsky [5, p. 69] generalized Saran’s $F_{K}$-function of three variables to $L$ variables, i.e.,
@@
Here it is worth mentioning that Khichi has introduced the multivariate hypergeometric functions $H_{B}^{(n)}$ ([49, p. 308, Eq. (123)]), which appears quite similar to the function (3.3), although they are different in nature.
@@ 4.1 A Joshi-Vyas type theorem
Inspired by the formula [22, p. 133] of Joshi and Vyas and its useful $q$-analogue [23, p. 646], we establish–using the measure in (21)–the following theorem of Joshi-Vyas type.
@@
When $k=1$, the formula (4.1) reduces to the result obtained by Joshi and Vyas. It can also be considered as Ernst’s $q$-analogue for Exton’s result (see [15, p. 189, Theorem 3.1]).
@@
This integral (4.3) is clearly a $q$-analogue of the integral by Luo et al. [33, Theorem 1]. In particular, when $x=0$, (4.3) reduces to
@@
which is the $q$-analogue of Koschmieder’s integral representation for $F_{2}$ ([33, Corollary 1]). Here, $\Phi_{2}$ and $\Phi_{0:2;2}^{1:2;2}$ denote the $q$-Appell function [20, p. 283, Eq. (10.2.6)] and the $q$-Kampé de Fériet function [49, p. 349], respectively.
@@
This integral is the $q$-analogue of the integral by Luo and Raina [32, p. 14, Theorem 4.1].
@@
Let $f:\mathbb{C}\rightarrow\mathbb{C}$ be a function of one variable. We define the $q$-shift operator $\mathbb{E}_{q}$ by $\mathbb{E}_{q}f(x)=f(qx)$ $(0<q<1)$ [30, Definition 1.3]. Inductively, for any $n\in\mathbb{Z}_{\geq 1}$,
@@
where $\left[\begin{matrix}k\\ p\end{matrix}\right]_{q}$ denotes the $q$-binomial coefficient [20, p. 24]. Then the use of the $q$-binomial theorem [20, p. 25] gives
@@
5 Conclusion
In this paper, we continue the recent work of Luo, Xu, and Raina [33] on Erdélyi-type integrals for Saran’s function $F_{K}$, providing a new proof of one of their main results (Theorem 1.1). Building on this, we obtain an important generalization, which is applicable to the $L$-variable $F_{K}$ function (3.3) that arises in physics. Furthermore, some interesting $q$-analogues and discrete analogues are also discussed.
By reviewing the literature on Erdélyi-type integrals of various hypergeometric functions over the past century, this paper compiles the table in Appendix 6, which reflects the overall landscape of research in this direction. As already clarified in Section 1, whether out of computational interest or due to their usefulness, it is highly valuable to continue exploring Erdélyi-type integrals for other hypergeometric functions not yet included in this table.
@@ 6 Appendix
others Volkodavov-Nikolaev (1993) [56]
Author Contributions
All the authors have contributed equally. All authors have read and approved the final manuscript.
@@
The research of the second author is supported by National Natural Science Foundation of China (Grant No. 12001095).
@@ Acknowledgements
Conflicts of interest
The authors declare that there is no conflict of interest.